# PSZ2 G181.06+48.47 I: X-ray exploration of a low-mass cluster with exceptionally-distant radio relics

### v1

@@ 7 The Merger Scenario
PSZ2 G181.06+48.47 is a low-mass post-core passage system, resulting from the merger of two subclusters along the N-NE–S-SW axis. The post-merger scenario is strongly supported by the bridge of cool, low-entropy gas connecting the two subcluster cores which indicates gas stripping during the first passage, suggesting a post-merger scenario. Although the dynamical mass is biased high, the X-ray, WL, and Planck mass estimates ($(2.5-4.2)\times 10^{14}\;\mathrm{M_{\odot}}$) confirm that PSZ2 G181.06+48.47 is one of the lowest-mass systems to host a double radio relic system.
While double relic systems are typically assumed to be less affected by projection effects, PSZ2 G181.06+48.47 might not conform to this expectation. While the X-ray data do not directly constrain the orientation of the merger in the plane of the sky, the symmetric nature, comparable sizes, and thin width of the two radio relics and the similar properties of the subclusters would, at first glance, suggest that the merger is occurring in the plane-of-the-sky. However, based on a radio polarization and color-color analysis, Rajpurohit et al. (submitted) ruled out the possibility that both relics are seen close to edge-on. Instead, Rajpurohit et al. (submitted) presented compelling evidence that the system could be tilted at an angle as high as $\sim 45^{\circ}$. The northern BCG is also located at a higher LOS velocity ($\Delta v\sim 980$ km s-1) compared to the southern one, lending further support that the merger axis is tilted with respect to the plane-of-the-sky. Takizawa et al. (2010) provided insights from hydrodynamical simulations indicating that the dynamical/virial mass can be more than twice the actual mass when observing along the collision axis at $t=1.33$ Gyr after the first core passage.
The $\sim 370$ kpc projected separation between the X-ray peaks of the two subclusters is dwarfed by comparison to the $\sim 2.6$ Mpc projected separation between the two radio relics, complicating the interpretation of how double relics can form at such large separation, where the density of particles available for acceleration is very low (e.g. Brunetti & Jones, 2014; Botteon et al., 2020b). As discussed above, if the merger axis is not in the plane of the sky, the two relics would be physically located even farther apart from each other ($\sim 3.7$ Mpc Rajpurohit et al., submitted). The limited time available for evolution would then present new challenges in explaining how relics propagate to such large distances. Rajpurohit et al. (submitted) propose that the relics are associated with shock re-acceleration of particles previously accelerated through, for example, active galactic nuclei, or re-acceleration through multiple sequential shocks. The relics follow typical scaling relations, having typical LLSs and powers, with their large cluster-centric being the main remarkable feature.
@@
If the subclusters are seen in the outgoing phase, after the second core passage, the small core distance would be easily understood. A post-apocenter scenario would naturally lead to X-ray edges close to the core, which could be formed more recently during the second encounter of the two cores, similar to the scenario proposed for CIG 0217+70 by Tümer et al. (2023). The inner shocks would then be driven by the second infall of the subclusters post-apocenter. However, compared to the post-apocenter scenario, the time added to cross the cores would likely lead to even larger separations between the relics, as discussed through simulations by Ahn et al. (in prep). Furthermore, the bridge would likely be completely destroyed and mixed in as the two subclusters pass through each other for a second time as indicated in simulations by ZuHone (2011).
@@ 8 Conclusion
We explored new Chandra and XMM-Newton X-ray observations of the low-mass merging cluster PSZ2 G181.06+48.47, which hosts two spectacular symmetrically located double relics. We reveal that the cluster is very cool ($kT_{500}=3.86^{+0.27}_{-0.29}$ keV) and has a mass, $M_{500,X}=2.57^{+0.37}_{-0.38}\times 10^{14}\;{\rm M_{\odot}}$, significantly lower than that previously measured by Planck, but consistent with the estimate from the WL (Ahn et al., in prep). Among the lowest mass clusters ever discovered to host double radio relics, the system is a major merger along the N-NE–S-SW direction, between two clearly-identified subclusters with a mass ratio 1.2–1.4.
@@
We discover a set of three, weak ($\mathcal{M}_{X,\rho}\sim 1.3-1.4$) shocks located within $<500$ kpc of the system center and aligned with the merger axis. The lack of significant density jumps provides an upper limit of $\mathcal{M}_{\rm X,\rho}\lesssim 1.6$ for the NE and SW relic shocks, significantly lower Mach numbers compared to estimations from the radio data (Rajpurohit et al., submitted), suggesting an inclination of the system with respect to the plane-of-the-sky.
Through a careful analysis of the existing literature, we compiled a sample of all known double radio relic systems and revised scaling relations by adding 12 new systems not included in previous work. We find positive correlations between relics power vs mass, relics power vs LLS, LLS vs relic-cluster distance, and LLS vs relic-relic distance. The radio relic power at 1.4 GHz scales as $P_{1.4\,\mathrm{GHz}}\propto M_{500}^{3.10\pm 0.59}$, consistent with previous findings. We find that the relics in PSZ2 G181.06+48.47 have typical luminosities and sizes, but they are located exceptionally far away from the cluster center, around $r_{200,WL}$, considering the low mass of the cluster. The nature of the diffuse radio sources in PSZ2 G181.06+48.47 is investigated in Rajpurohit et al. (submitted).
The existing observational evidence points to a late-stage, post-merger scenario with a significant LOS component, where the two subclusters are seen post-apocenter as they are falling back into each other. The outer shocks produced during the first encounter expand outwards, reaching large radii over time, which would account for the large distance between the two relics. The merger scenario is further explored through simulations in Ahn et al. (in prep).
@@ Acknowledgments
We thank Felipe Andrade-Santos, Tim Shimwell, Grant Tremblay and Christine Jones for useful discussions. A. Stroe, Z. Zhu, and K. Rajpurohit acknowledge support from the NASA 80NSSC21K0822 and Chandra GO0-21122X grants. A. Stroe gratefully acknowledges previous support from a Clay Fellowship. Z. Zhu and A. Simionescu are supported by the Netherlands Organisation for Scientific Research (NWO). L. Lovisari acknowledges support from INAF grant 1.05.12.04.01. W. Forman acknowledges support from the Smithsonian Institution, the Chandra High Resolution Camera Project through NASA contract NAS8-03060, and NASA Grants 80NSSC19K0116 and GO1-22132X. M. James Jee acknowledges support for the current research from the National Research Foundation (NRF) of Korea under the programs 2022R1A2C1003130 and RS-2023-00219959. This work is based on observations obtained with XMM-Newton, an ESA science mission with instruments and contributions directly funded by ESA Member States and the US (NASA). This research has made use of observations made by the Chandra X-ray Observatory, and software provided by the Chandra X-ray Center (CXC). LOFAR (van Haarlem et al., 2013) is the Low Frequency Array designed and constructed by ASTRON.
Appendix A Masked point sources
Figure 12: XMM-Newton flux image showing in pink ellipses the point sources and extended features excluded from the analysis. The two clumps associated with background clusters are marked in cyan.
Appendix B Temperature uncertainty map
@@
Appendix C $\log N-\log S$ plot
Figure 14: $\log N-\log S$ curve of XMM-Newton-detected point sources in PSZ2 G181.06+48.47, with the best-fit power-law model overplotted. The curve from CDF-S is plotted (Lehmer et al., 2012). $1\sigma$ uncertainty ranges are denoted in the shaded regions. The XMM-Newton detection limit, which was used as the point source exclusion flux limit in the spectral analysis, is also shown. Only sources located within $3-10\arcmin$ of the cluster center were included.
Appendix D Galaxy clusters hosting double radio relics
In Table LABEL:table:double_radio_relic, we present the properties of our compilation of double radio relics in galaxy clusters known to date. For relics lacking 1.4 GHz flux density measurements, we either assumed a radio spectral index of $\alpha=-1.3$ to convert to the equivalent 1.4 GHz radio power or extrapolated the 1.4 GHz flux density based on their low-.frequency spectrum

### v3

@@ 7 The Merger Scenario
PSZ2 G181.06+48.47 is a low-mass post-core passage system, resulting from the merger of two subclusters along the N-NE–S-SW axis. The post-merger scenario is strongly supported by the bridge of cool, low-entropy gas connecting the two subcluster cores which indicates gas stripping during the first passage. Although the dynamical mass is biased high, the X-ray, WL, and Planck mass estimates ($(2.5-4.2)\times 10^{14}\;\mathrm{M_{\odot}}$) confirm that PSZ2 G181.06+48.47 is one of the lowest-mass systems to host a double radio relic system.
While double relic systems are typically assumed to be less affected by projection effects, PSZ2 G181.06+48.47 might not conform to this expectation. While the X-ray data do not directly constrain the orientation of the merger in the plane-of-the-sky, the symmetric nature, comparable sizes, and thin width of the two radio relics and the similar properties of the subclusters would, at first glance, suggest that the merger is occurring in the plane-of-the-sky. However, based on a radio polarization and color-color analysis, Rajpurohit et al. (2025) ruled out the possibility that both relics are seen close to edge-on. Instead, Rajpurohit et al. (2025) presented compelling evidence that the system could be tilted at an angle as high as $\sim 45^{\circ}$. The northern BCG is also located at a higher LOS velocity ($\Delta v\sim 980$ km s-1) compared to the southern one, lending further support that the merger axis is tilted with respect to the plane-of-the-sky. Takizawa et al. (2010) provided insights from hydrodynamical simulations indicating that the dynamical/virial mass can be more than twice the actual mass when observing along the collision axis at $t=1.33$ Gyr after the first core passage. The possibility of a significant impact parameter, especially along the LOS, is further discussed later on.
The $\sim 370$ kpc projected separation between the X-ray peaks of the two subclusters is dwarfed by comparison to the $\sim 2.6$ Mpc projected separation between the two radio relics, complicating the interpretation of how double relics can form at such large separation, where the density of particles available for acceleration is very low (e.g. Brunetti & Jones, 2014; Botteon et al., 2020b). As discussed above, if the merger axis is not in the plane-of-the-sky, the two relics would be physically located even farther apart from each other ($\sim 3.7$ Mpc Rajpurohit et al., 2025). The limited time available for evolution would then present new challenges in explaining how relics propagate to such large distances. Rajpurohit et al. (2025) propose that the relics are associated with shock re-acceleration of particles previously accelerated through, for example, active galactic nuclei, or re-acceleration through multiple sequential shocks. The relics follow typical scaling relations, having typical LLSs and powers, with their large cluster-centric being the main remarkable feature.
@@
If the subclusters are seen in the outgoing phase, after the second core passage, the small core distance would be easily understood. A post-apocenter scenario would naturally lead to X-ray edges close to the core, which could be formed more recently during the second encounter of the two cores, similar to the scenario proposed for CIG 0217+70 by Tümer et al. (2023). The inner shocks would then be driven by the second infall of the subclusters post-apocenter. However, compared to the post-apocenter scenario, the time added to cross the cores would likely lead to even larger separations between the relics, as discussed through simulations by Ahn et al. (2025). Furthermore, the bridge would likely be completely destroyed and mixed in as the two subclusters pass through each other for a second time as indicated in simulations by ZuHone (2011).
@@ 8 Conclusion
We explored new Chandra and XMM-Newton X-ray observations of the low-mass merging cluster PSZ2 G181.06+48.47, which hosts two spectacular symmetrically located double relics. We reveal that the cluster is very cool ($kT_{500}=3.62^{+0.15}_{-0.07}$ keV) and has a mass, $M_{500,X}=2.32^{+0.29}_{-0.25}\times 10^{14}\;{\rm M_{\odot}}$, significantly lower than that previously measured by Planck, but consistent with the estimate from the WL (Ahn et al., 2025). Among the lowest mass clusters ever discovered to host double radio relics, the system is a major merger along the N-NE–S-SW direction, between two clearly-identified subclusters with a mass ratio 1.2–1.4.
@@
We discover a set of three, weak ($\mathcal{M}_{X,\rho}\sim 1.3-1.4$) shocks located within $<500$ kpc of the system center and aligned with the merger axis. The lack of significant density jumps provides an upper limit of $\mathcal{M}_{\rm X,\rho}\lesssim 1.6$ for the NE and SW relic shocks, significantly lower Mach numbers compared to estimations from the radio data (Rajpurohit et al., 2025), suggesting an inclination of the system with respect to the plane-of-the-sky.
Through a careful analysis of the existing literature, we compiled a sample of all known double radio relic systems and revised scaling relations by adding 12 new systems not included in previous work. We find positive correlations between relics power vs mass, relics power vs LLS, LLS vs relic-cluster distance, and LLS vs relic-relic distance. The radio relic power at 1.4 GHz scales as $P_{1.4\,\mathrm{GHz}}\propto M_{500}^{3.10\pm 0.59}$, consistent with previous findings. We find that the relics in PSZ2 G181.06+48.47 have typical luminosities and sizes, but they are located exceptionally far away from the cluster center, around $r_{200,WL}$, considering the low mass of the cluster. The nature of the diffuse radio sources in PSZ2 G181.06+48.47 is investigated in Rajpurohit et al. (2025).
The existing observational evidence points to a late-stage, post-merger scenario with a significant LOS component, where the two subclusters are seen post-apocenter as they are falling back into each other. The outer shocks produced during the first encounter expand outwards, reaching large radii over time, which would account for the large distance between the two relics. The merger scenario is further explored through simulations in Ahn et al. (2025).
@@ Acknowledgments
We thank the anonymous referee for their useful feedback. We thank Felipe Andrade-Santos, Tim Shimwell and Grant Tremblay for useful discussions. A. Stroe, Z. Zhu, and K. Rajpurohit acknowledge support from the NASA 80NSSC21K0822 and Chandra GO0-21122X grants. A. Stroe gratefully acknowledges previous support from a Clay Fellowship. Z. Zhu and A. Simionescu are supported by the Netherlands Organisation for Scientific Research (NWO). L. Lovisari acknowledges support from INAF grant 1.05.12.04.01. W. Forman acknowledges support from the Smithsonian Institution, the Chandra High Resolution Camera Project through NASA contract NAS8-03060, and NASA Grants 80NSSC19K0116 and GO1-22132X. M. James Jee acknowledges support for the current research from the National Research Foundation (NRF) of Korea under the programs 2022R1A2C1003130 and RS-2023-00219959. This work is based on observations obtained with XMM-Newton, an ESA science mission with instruments and contributions directly funded by ESA Member States and the US (NASA). This research has made use of observations made by the Chandra X-ray Observatory, and software provided by the Chandra X-ray Center (CXC). LOFAR (van Haarlem et al., 2013) is the Low Frequency Array designed and constructed by ASTRON.
Appendix A Temperature uncertainty map
@@
Appendix B Galaxy clusters hosting double radio relics